The wall of a reservoir on which the slide gate is located is inclined at 300 to the horizontal.
The gate is hinged at its upper end and a massive weight is pushed against its lower end to keep it closed.
Determine the reaction of the weight and its direction if the gate has a circular cross
section of diameter 3m.
1
Expert's answer
2021-08-30T01:54:10-0400

"Reaction = \gamma Ah \cos \theta\\
Reaction =9790* \frac{\pi}{4}(3)^2*1* \cos 30\\
Reaction=3964950^{\circ \:}\cdot \frac{\sqrt{3}}{2}\\
Reaction=59930.19824N"